# High-speed ditching of double curvature specimens with cavitation and ventilation

### v1

@@ Acknowledgements
The authors also wish to thank: Flavio Olivieri, the technical supervisor of the High Speed Ditching Facility, who is also responsible for its maintenance. He also participated in the the execution of all the experiments; Fabio Carta, the project manager who dealt with the design of the specimens and with the installation of the instrumentation on board of them and on the facility; Mauro Sale who provided technical support during the execution of the experiments; Alessandro Del Buono was of value in providing assistance with the interpretation of the results.
@@
This project has been partly funded from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No. 724139 (H2020-SARAH: increased SAfety & Robust certification for ditching of Aircrafts & Helicopters).
@@
Author contributions
Alessandro Iafrati and Silvano Grizzi designed the facility, installed the experimental setup and performed the experiments. Emanuele Spinosa and Silvano Grizzi carried out the data analysis. Emanuele Spinosa organized and prepared the article drafts. Alessandro Iafrati revised the drafts and supervised the article preparation.

### v2

@@ Acknowledgements
The authors also wish to thank: Flavio Olivieri, the technical supervisor of the High Speed Ditching Facility, who is also responsible for its maintenance. He also participated in the the execution of all the experiments; Fabio Carta, the project manager who dealt with the design of the specimens and with the installation of the instrumentation on board of them and on the facility; Mauro Sale who provided technical support during the execution of the experiments; Alessandro Del Buono who gave valuable support in the interpretation of the results.
@@
This project has been funded from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No. 724139 (H2020-SARAH: increased SAfety & Robust certification for ditching of Aircrafts & Helicopters).